Lalasa and Yasmin are designing a triangular banner to hang in the school gymnasium. They first draw the design on paper. The triangle has a base of 5 inches and a height of 7 inches. If 1 inch on the drawing is equivalent to 1.5 feet on the actual banner, what will the area of the actual banner be?
Answer by Theo(13342)

the triangle on paper is 5 inches wide by 7 inches high.
if 1 inch on paper is equal to 1.5 feet on paper, then you need to multiply all dimensions by 18 because 1.5 feet is equal to 18 inches and you are getting 18 inches on the real banner for every inch on the paper banner.

5 * 18 = 90 inches for the bsse of the real triangle.
7 * 18 = 125 inches for the height of the real triangle.

the area of the paper triangle will be 5 * 7 / 2 = 35/2 square inches.
the area of the real triangle will be 90 * 125 / 2 = 5625 square inches.

in terms of feet:

the base of the real banner will be 125/12 = 10.41666666....... feet.
the height of the real banner will be 90 / 12 = 7.5 feet.
the area of the real banner will be 5625 / 12^1 = 5625 / 144 = 39.0625 square feet.